10.00 points Suppose your irm has decided to use a divisional WACC approach to analyze projects. The firm currently has four divisions, A through D, with average betas for each division of 0.8, 1.2, 1.4, and 1.6, respectively Assume all current and future projects will be financed with 20 percent debt and 80 percent equity, the current cost of equity (based on an average firm beta of 1 1 and a current risk-free rate of 7 percent) is 11 percent and the after-tax yield on the companys bonds is 12 percent What will the WACCs be for each division? (Do not round intermediate calculations and round your final answers to 2 decimal places.) WACCS Division A Division B Division C Division D
Answer #1

1st we need to calculate the Rm (Return of market.)

Given, current cost of equity = 11% with beta of 1.1 and Rf = 7%.

As per CAPM we know Re=Rf + B(Rm-Rf)

or, 11% = 7% + 1.1(Rm - 7%)

or, Rm = {(11 - 7) / 1.1}+7

or, Rm = 10.64%.

DIVISIONAL A

Return on equity (Re) = 7% + 0.8(10.64 - 7)% i.e. 9.91%

Return on Debt (Rd) = 12%

WACC = (Re * We) + (Rd * Wd)

WACC = (9.91% * 0.8) + (12% *0.20)

WACC = 10.33%

DIVISIONAL B

Return on equity (Re) = 7% + 1.2(10.64 - 7)% i.e. 11.37%

Return on Debt (Rd) = 12%

WACC = (Re * We) + (Rd * Wd)

WACC = (11.37% * 0.8) + (12% *0.20)

WACC = 11.50%

DIVISIONAL C

Return on equity (Re) = 7% + 1.4(10.64 - 7)% i.e. 12.10%

Return on Debt (Rd) = 12%

WACC = (Re * We) + (Rd * Wd)

WACC = (12.10% * 0.8) + (12% *0.20)

WACC = 12.08%

DIVISIONAL D

Return on equity (Re) = 7% + 1.6(10.64 - 7)% i.e. 12.82%

Return on Debt (Rd) = 12%

WACC = (Re * We) + (Rd * Wd)

WACC = (12.82% * 0.8) + (12% *0.20)

WACC = 12.66%